OpenStudy (anonymous):

in a sample of 50 museum visitors in one afternoon, 22 were female and 28 were male. if there were 800 museum visitors that afternoon, estimate the number of visitors who were female a. 220 b. 352 c. 448 d. 700

OpenStudy (anonymous):

352 (22/50).800 = 22*16 = 352

OpenStudy (anonymous):

thank you! :D

OpenStudy (anonymous):

You're Welcome!:)